OpenStudy (anonymous):

Solve. (x + 2)2 = 1 x = –3 and x = –1 x = 3 and x = 1 x = –1 and x = 4 x = –4 and x = –3

OpenStudy (theviper):

\[(x+2)^2=1\]\[x^2+4x+4=1\]\[x^2+4x+4-1=0\]\[x^2+4x+3=0\]

OpenStudy (anonymous):

ok now what

OpenStudy (theviper):

\[x^2+x+3x+3=0\]\[x(x+1)+3(x+1)=0\]\[(x+3)(x+1)=0\] x=-3 or x=-1

OpenStudy (theviper):

understood??

OpenStudy (anonymous):

@harrisonshields simply get the square root of both sides as such: \[\sqrt{(x+2)^{2}} = \sqrt{1}\] \[x+2 = \pm 1\] so from here x+2 =1+or x+2 = -1 can you proceed from here?